OpenStudy (anonymous):

If x>1, which of the following increase(s) as x increases? 1. x-1/x II. 1/x^2-x III. 4x^3-2x^2 (A) Only I (B) Only II (C) Only III (D) Only I and III (E) All

terenzreignz (terenzreignz):

Is this a calculus question? :D

OpenStudy (anonymous):

Umm...I think soo..Not sure..

terenzreignz (terenzreignz):

Derivatives? :)

OpenStudy (anonymous):

Dont know what that is...

terenzreignz (terenzreignz):

We have a problem then... lol let's see..

terenzreignz (terenzreignz):

What's the second one, by the way? \[\Large \frac1{x^2}-x\]? :D

OpenStudy (anonymous):

Close. ^^ But the "-x" goes on the denominator..

terenzreignz (terenzreignz):

Close is just a fancy word for 'wrong' XD \[\Large \frac1{x^2 - x}\]

terenzreignz (terenzreignz):

Okay, so let's look at the first one, and analyse a bit... \[\Large \frac{x-1}{x}\] look good? :D

OpenStudy (anonymous):

Yes =p

terenzreignz (terenzreignz):

Right, this can be simplified into \[\Large \frac{x-1}x = \frac{x}x - \frac1x = 1 - \frac1x\] catch me so far?

OpenStudy (anonymous):

noo...I get the first and the last one but not the middle one..

terenzreignz (terenzreignz):

Well... \[\Large \frac{x-1}x= \frac1x(x - 1) = \frac{x}x-\frac1x\] :)

OpenStudy (anonymous):

Ohhhh...........

OpenStudy (anonymous):

What next?^^

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

Could you clarify what the functions are? Maybe draw them? I'm not sure what goes in the numerator and what goes in the denominator.

OpenStudy (anonymous):



OpenStudy (anonymous):